OpenStudy (anonymous):

A remote-control car races straight down the street at 15 feet per second. Four seconds later, a second remote-control car races straight down the same street at 25 feet per second. How long will it take the second car to catch up to the first car? A remote-control car races straight down the street at 15 feet per second. Four seconds later, a second remote-control car races straight down the same street at 25 feet per second. How long will it take the second car to catch up to the first car? @Mathematics

OpenStudy (anonymous):

15 x 4 = 60 feet (total head start) 60 / 25 = 2.4 sec (total head start / 2nd car speed) 15 x 2.4 = 36 feet (additional feet gained by 1st car while 2nd car makes up 60 feet) 36 + 60 = 96 feet (additonal feet + headstart) 96 / 25 = 3.84 sec (Total feet / 2nd car speed) 3.84 + 4 = 7.84 (additional time + headstart) My answer is 7.84 sec

OpenStudy (anonymous):

Solve as an equation, where distance travelled by first car=distance travelled by second car: 15*4+ 15x=25x (Where x is the number of seconds after the second car starts). 10x=60, which gives x=6 seconds